# A Survey of Behavior Foundation Model: Next-Generation Whole-Body Control System of Humanoid Robots

### v4

@@ I Introduction
Humanoid robots are increasingly being developed and deployed in a variety of real-world scenarios due to their human-like morphology and high degrees of freedom (DoF) [darvish2023teleoperation, valenzuela2024embodying, tong2024advancements]. These capabilities allow them to operate seamlessly in environments originally designed for humans, allowing them to perform locomotion, manipulation, and interaction tasks with versatility and agility. To enable such rich behaviors, whole-body control (WBC) plays a central role by coordinating numerous joints, contacts, and actuators to achieve balanced, agile, and safe full-body motions. However, designing effective WBC policies remains highly challenging due to the complex dynamics of humanoids, underactuation, frequent contact changes, and the strong coupling between locomotion, manipulation, and balance [kuindersma2016optimization, cheng2024expressive]. These difficulties highlight the need for more general, robust, and scalable control frameworks. In the following sections, we begin by comprehensively reviewing previous humanoid WBC methods, ranging from traditional model-based to learning-based and task-specific controllers, before introducing the transformative approach—the behavior foundation model. This evolutionary trajectory not only reflects the field’s advancement toward enhanced intelligence and generalizability but also paves the way for the next-generation humanoid robot control systems.
@@ I-A Traditional Model-based Controller
Traditional WBC methods have served as the cornerstone for locomotion and manipulation in early humanoid robots [kulic2016anthropomorphic, goswami2019humanoid]. These methods rely heavily on physics-based models and are typically structured into a predictive-reactive hierarchy: high-level planners like centroidal model predictive control (MPC) [schwenzer2021review, romualdi2022online] generate reference trajectories, while low-level task-space whole-body controllers solve optimal control problems (OCPs) [sethi2021optimal] to track these objectives under dynamic constraints. For instance, operational space control [khatib2003unified] and hierarchical task control [sentis2005synthesis] establish the theoretical foundation, while [escande2014hierarchical] advances hierarchical quadratic programming (QP) solvers and enables real-time performance in multi-task scenarios such as balance, walking, and manipulation. Whole-body operational space control (WBOSC) extends these methods by integrating the control of multiple body parts, handling redundancy, and managing dynamic constraints across the entire robot body [sentis2006whole, kim2016stabilizing, fok2016controlit]. These frameworks have been widely applied in humanoids, such as Atlas [kuindersma2016optimization], HRP-2 [10000129], and DLR’s torque-controlled robots [henze2015approach], achieving robust locomotion and multi-contact interactions.
Despite their success, traditional WBC systems face critical limitations: (i) task design, gain tuning, and heuristic adjustments for complex behaviors (e.g., uneven terrain or dynamic transitions) remain labor-intensive and brittle, (ii) real-time MPC struggles with high-dimensional systems, often requiring simplifications that sacrifice dynamic fidelity [sovukluk2023whole], (iii) the lack of flexibility to execute highly dynamic skills (e.g., backflips or rapid contact switches) or adapt to unforeseen disturbances [ishihara2019full], and (iv) weak robustness, as even a little push may topple a robot with model-based walking controller. These challenges are especially significant in humanoids, where tasks often require rich coordination, contact reasoning, and situational awareness [hao2025embodied, murooka2024whole]. As a result, recent research increasingly shifts toward data-driven approaches, aiming to learn motor skills, coordination policies, and behavioral priors from demonstrations or reinforcement learning [peng2018deepmimic, singh2021parrot].
@@ I-B Learning-based and Task-specific Controller
Learning-based methods, particularly reinforcement learning (RL) and imitation learning (IL), have emerged as promising alternatives to traditional WBC methods, enabling robots to acquire complex skills through environmental interaction or human demonstrations [kober2013reinforcement, sunderhauf2018limits, nguyen2019review, karoly2020deep, liu2021deep, huang2022reward, zhang2022adjacency, wu2023human, chen2023hierarchical, zhu2023transfer, luo2024moil]. For example, the DeepMimic framework [peng2018deepmimic] combines deep RL with motion capture data, allowing physically simulated characters to learn dynamic skills while maintaining natural motion quality. This work is further extended by adversarial motion priors (AMP) [peng2021amp], which enables more stylized and diverse character control while preserving physical realism. Similarly, the HoST [huang2025learning] framework introduces an RL-based standing-up controller that learns adaptive and stable motions across diverse laboratory and outdoor environments, highlighting the robustness and task-specific learning capability of RL. For IL-based approaches, the TRILL [seo2023deep] approach combines virtual reality (VR) teleoperation with WBC for humanoid loco-manipulation, demonstrating an 85% success rate in real-world bimanual tasks [bang2023control]. Further advances, such as ExBody [cheng2024expressive, ji2025exbody2], develop an expressive WBC framework that decouples upper-body IL (for stylistic motions) from robust lower-body locomotion, enabling humanoid robots to dynamically adapt their gait while performing diverse movements. This approach overcomes the instability of full-body imitation caused by morphological mismatches between humans and robots.
While learning-based methods have demonstrated remarkable success in diverse humanoid WBC tasks, they face fundamental challenges that limit their broader applicability. RL-based approaches suffer from sample inefficiency, often requiring millions of environment interactions to converge, while remaining highly sensitive to the reward function design—poorly shaped rewards can lead to unintended behaviors or local optima [maeureka]. Furthermore, the simulation-to-reality (Sim2Real) gap exacerbates these limitations, as policies trained in simulation frequently degrade when confronted with real-world dynamics, sensor noise, and hardware imperfections [kadian2020sim2real, hofer2021sim2real, iyengar2023sim2real, su2024sim2real]. In contrast, IL-based methods are more sample-efficient yet pose a significant challenge to data collection, and learned policies often inherit the biases and limitations of the demonstrator [schaal1999imitation, hussein2017imitation, laskey2017dart, fang2019survey, hua2021learning]. Moreover, both paradigms struggle with generalization, in which learned policies typically excel only at narrow tasks and fail to adapt to new scenarios without extensive retraining. These challenges collectively underscore the need for approaches that combine the flexibility of learning with structured priors for robustness and generalizability, a gap that behavior foundation models aim to bridge.
@@ I-C Behavior Foundation Model
The term ”behavior(al) foundation model (BFM)” is first introduced in [pirotta2024fast], which proposes a successor measure-based framework for training generalist policies capable of instantly imitating diverse behaviors from minimal demonstrations. It demonstrates that BFMs pretrained on unsupervised interaction data are promising for eliminating task-specific RL fine-tuning by solving imitation tasks through forward-backward state feature matching, while simultaneously supporting multiple IL paradigms via a unified representation, such as behavioral cloning, reward inference, and distribution matching. Subsequent work [cetin2025finer, tirinzoni2025zeroshot, sikchi2025fast, vainshtein2025task, bobrin2025zero] has established BFMs as a class of RL agents capable of unsupervised training on reward-free transitions while yielding approximately optimal policies for broad classes of reward functions at test time without additional learning or planning.
In this paper, we extend the definition of BFMs as a specialized class of foundation models [bommasani2021opportunities] designed to control agent behaviors in dynamic environments. As illustrated in Fig. 1, rooted in the principles of general foundation models (e.g., GPT-4 [achiam2023gpt], CLIP [radford2021learning], and SAM [kirillov2023segment]) that leverage broad, self-supervised pre-training on large-scale static data, BFMs are often trained on extensive behavior data (e.g., human demonstrations, or agent-environment interactions), encoding a comprehensive spectrum of behaviors rather than specializing narrowly in single-task scenarios. This property ensures that the model can readily generalize across different tasks, contexts, or environments, demonstrating versatile and adaptive behavior generation capabilities. Recent advancements in vision-language-action (VLA) models [radford2021learning, zitkovich2023rt, kim2024openvla, wen2025tinyvla] have focused on integrating vision, language, and action to handle multi-modal tasks, excelling in dynamic settings where they generate context-aware responses based on visual and linguistic inputs. In contrast, BFMs are primarily designed for directly controlling agent behaviors such as locomotion, manipulation, and interaction. Moreover, most existing VLA models are applicable to relatively stable platforms, such as mechanical arms or wheeled humanoid robots [zambella2019dynamic], whereas BFMs are designed to handle the sophisticated WBC of full-size humanoid robots.
@@ II-A Humanoid Whole-body Control
Humanoid robots are expected to operate in unstructured environments that are dynamic and unpredictable, demanding control systems that are highly versatile, robust, reconfigurable, dexterous, and mobile compared to less agile robotic platforms [moro2019whole]. To that end, the humanoid WBC is proposed to coordinate the motion of multiple robots’ appendages to execute multiple tasks simultaneously and reliably. It considers the entire robot body as a single and integrated system, managing locomotion, manipulation, and interaction with the environment using a unified set of control algorithms [sentis2006whole]. As illustrated in Fig. 2, the humanoid WBC has evolved from traditional model-based approaches to flexible learning-based approaches, moving toward a generalist that solves broad tasks in diverse scenarios [moro2019whole, zambella2019dynamic, cheng2024expressive, xue2025unified, he2025hover, wang2025experts]. In line with this trend, BFMs have emerged as a promising approach to achieve general-purpose WBC through large-scale pre-training on diverse motion data, which will be comprehensively discussed in the following contents.
@@ II-B Reinforcement Learning
We frame the RL problem as a Markov decision process (MDP), which can be defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r^{\rm ext},\gamma)$ [sutton1998reinforcement]. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(\bm{s}^{\prime}|\bm{s},\bm{a})$ is the probability measure on $\mathcal{S}$ defining the stochastic transition to the next state $\bm{s}^{\prime}$ obtained by taking action $\bm{a}$ in state $\bm{s}$, $r^{\rm ext}(\bm{s}):\mathcal{S}\rightarrow\mathbb{R}$ is the extrinsic/task reward function, and $\gamma\in[0,1]$ is a discount factor. A policy $\pi$ is defined as the probability measure $\pi(\bm{a}|\bm{s})$ that maps each state to a distribution over actions. Furthermore, we denote $\mathrm{Pr}(\cdot|\bm{s}_{0},\bm{a}_{0},\pi)$ and $\mathbb{E}(\cdot|\bm{s}_{0},\bm{a}_{0},\pi)$ probability and expectation operators under state-action sequences $(\bm{s}_{t},\bm{a}_{t})_{t\geq 0}$ starting at $(\bm{s}_{0},\bm{a}_{0})$ and following the policy $\pi$ with $\bm{s}_{t}\sim P(\bm{s}_{t}|\bm{s}_{t-1},\bm{a}_{t-1})$ and $\bm{a}_{t}\sim\pi(\bm{a}_{t}|\bm{s}_{t})$. The goal of RL is to learn a policy that maximizes the expected discounted return:
@@ III-A1 Goal-conditioned Learning
As shown in Figure 4, goal-conditioned learning is a framework in RL where an agent’s behavior is conditioned on a specific goal or objective, typically provided as input. Unlike traditional RL, where the agent learns from raw state-action pairs without explicit task-specific guidance, goal-conditioned learning integrates the goal into the agent’s policy, enabling it to adapt its actions toward achieving that specific goal. The goal can be specified in various forms, such as a target state, an objective function, or an external task description [luo2023perpetual, wu2023masked, jiang2024hgap]. The key advantage of goal-conditioned learning lies in its ability to learn a more flexible and transferable policy that can be applied to a wide range of tasks, as it directly incorporates the task’s goal during training, rather than requiring retraining for each specific task. This makes it particularly useful in environments where the agent needs to solve multiple tasks or interact with changing environments.
@@
TeamPlay [liu2022motor], ASE [peng2022ase], PHC [luo2023perpetual], CALM [tessler2023calm], MTM [wu2023masked],
CASE [dou2023case], InterMimic [xu2025intermimic], MoConVQ [yao2024moconvq], H-GAP [jiang2024hgap],
MaskedMimic [tessler2024maskedmimic], HugWBC [xue2025unified], HOVER [he2025hover], ModSkill [huang2025modskill],
CLONE [li2025clone], TWIST [ze2025twist], TWIST2 [ze2025twist2], AMS [pan2025agility],
Any2Track [zhang2025track], BFM4Humanoid [zeng2025behavior], SONIC [luo2025sonic]
@@
ICM [pathak2017curiosity], DIAYN [eysenbach2018diversity], RND [burda2018exploration], APS [liu2021aps],
ProtoRL [yarats2021reinforcement], RE3 [seo2021state], ODPP [chen2023unified]
@@
FB [touati2021learning], FB-IL [pirotta2024fast], FB-AW [cetin2025finer], FB-ARE [cetin2025finer],
FB-AWARE [cetin2025finer], Motivo [tirinzoni2025zeroshot], BFM-Zero [li2025bfm]
Adaptation
Adaptation
@@
TokenHSI [pan2025tokenhsi], Belief-FB [bobrin2025zero], Rotation-FB [bobrin2025zero],
TaskTokens [vainshtein2025task], ReLA [sikchi2025fast], LoLA [sikchi2025fast]
@@
UniHSI [xiao2024unified], TokenHSI [pan2025tokenhsi], UniPhys [wu2025uniphys], CLoSD [tevet2025closing],
LangWBC [shao2025langwbc], LeVERB [xue2025leverb], BeyondMimic [liao2025beyondmimic], SENTINEL [wang2025sentinel]
Skill learning from motion tracking. Among the diverse approaches to goal-conditioned learning, tracking-based learning represents a specialized form where the target behavior is explicitly defined by dense reference supervision or guidance, typically derived from motion capture data or expert demonstrations. At each time step, the agent is often trained to track the given reference motion’s joint angle or kinematic pose of the next time step [peng2018deepmimic]. The primary motivation behind tracking-based learning is that learning to track a single pose is more achievable and general than directly imitating a whole motion, especially a complex motion.
For example, the TeamPlay [liu2022motor] framework trains an agent to imitate a large amount of football motion capture data via a DeepMimic-like [peng2018deepmimic] approach, aiming to learn a complete behavior set required for the football game. Then the agent is leveraged to sample substantial state-action pairs to train a neural probabilistic motor primitive (NPMP) model [merel2018neural] and derive a low-level latent-conditioned controller. Finally, the controller is applied for further drill learning by conducting RL with a drill-specific (e.g., follow, dribble, shoot, and kick-to-target) reward function. Here, the learned low-level controller can be viewed as a BFM as it learns realistic human-like movements based on motion capture data and can be rapidly adapted to diverse higher-level drill learning.
Similarly, the adversarial skill embeddings (ASE) framework [peng2022ase] learns a reusable latent space of motor skills by combining adversarial IL with unsupervised RL. Trained on unstructured motion data, ASE produces a latent-conditioned low-level controller capable of generating diverse and physically plausible behaviors, serving as a general-purpose motor prior to downstream tasks. Building on ASE, the conditional adversarial latent models (CALM) framework [tessler2023calm] incorporates a conditional discriminator to enable fine-grained control over generated motions through latent manipulation. This line of work is further extended by CASE [dou2023case], which introduces skill-conditioned IL with training techniques such as focal skill sampling and skeletal residual forces to enhance agility and motion diversity.
@@
While the methods above achieve efficient skill acquisition from large behavior datasets, the HugWBC [xue2025unified] method explores learning versatile locomotion skills without relying on pre-collected motion data. The framework automatically generates adaptive behaviors through a structured RL process, where a general command space dynamically produces feasible velocity, gait, and posture targets during training. By reformulating WBC as a self-supervised command-tracking problem, the work establishes a new direction for developing general-purpose humanoid controllers that learn robust skills through environmental interaction rather than data imitation.
@@
Moving beyond body-level skill learning, ModSkill [huang2025modskill] introduces a modular framework that decouples full-body motion into part-specific skills for individual body parts. This modularization allows for efficient and scalable learning, as each body part is controlled independently by a low-level controller driven by part-specific skill embeddings. ModSkill’s ability to focus on body-part-level skills makes it a powerful system for controlling complex motions and adapting learned behaviors across different tasks. By utilizing a skill modularization attention layer, ModSkill enhances the generalization of motor skills across various tasks, such as reaching or striking, thereby further improving task-specific adaptation.
@@
From primitive skills to high-level goal execution. The success of BFMs in learning diverse primitive skills has propelled the development of more advanced BFMs capable of interpreting and executing high-level goals, including language instructions and multi-task objectives. A notable example is MoConVQ [yao2024moconvq], which introduces a unified motion control framework based on discrete latent codes learned via vector quantized variational autoencoder (VQ-VAE). The model supports a wide range of downstream tasks, including motion tracking, interactive control, and text-to-motion generation, by offering a compact and modular representation. MoConVQ also integrates with large language models (LLMs) and enables the simulated agents to be directed via in-context language prompts, thereby bridging symbolic reasoning and physical control.
@@
Figure 5: MaskedMimic [tessler2024maskedmimic] is a representative BFM trained via goal-conditioned learning, which supports diverse control modalities and achieves seamless transitions between distinct tasks.
@@
Meanwhile, MaskedMimic [tessler2024maskedmimic] addresses physics-based character control as a general motion inpainting problem, producing full-body motions from partial descriptio, such as masked keyframes, objects, or text instructions. MaskedMimic involves a two-phase training process. First, a fully-constrained motion tracking controller is trained to imitate diverse reference motions. Then, a partially-constrained VAE-based policy distills this knowledge through masked goal conditioning. As a result, MaskedMimic can dynamically adapt to complex scenes and support applications ranging from VR control to complex human-object interaction (HOI), as illustrated in Fig. 5. In addition, InterMimic [xu2025intermimic] focuses on the HOI scenario and designs a two-stage teacher-student framework that distills imperfect motion capture interaction data into robust and physics-based controllers. Teacher policies are trained on subsets of noisy data and refined through simulation, then distilled into a student policy with RL-based fine-tuning. This curriculum strategy enables generalization across diverse interactions with high physical fidelity.
@@
For real-world robotic applications, HOVER [he2025hover] introduces a multi-mode policy distillation framework that enables humanoid robots to switch seamlessly between tasks (e.g., locomotion, manipulation, and navigation) using a single unified policy distilled from an oracle. This eliminates the need for task-specific controllers, demonstrating general-purpose control in real-world environments, akin to the versatility seen in MaskedMimic for virtual characters.

### v5

@@ I Introduction
Humanoid robots are increasingly being developed and deployed in a variety of real-world scenarios due to their human-like morphology and high degrees of freedom (DoF) [1, 2, 3]. These capabilities allow them to operate seamlessly in environments originally designed for humans, allowing them to perform locomotion, manipulation, and interaction tasks with versatility and agility. To enable such rich behaviors, whole-body control (WBC) plays a central role by coordinating numerous joints, contacts, and actuators to achieve balanced, agile, and safe full-body motions. However, designing effective WBC policies remains highly challenging due to the complex dynamics of humanoids, underactuation, frequent contact changes, and the strong coupling between locomotion, manipulation, and balance [4, 5]. These difficulties highlight the need for more general, robust, and scalable control frameworks. In the following sections, we begin by comprehensively reviewing previous humanoid WBC methods, ranging from traditional model-based to learning-based and task-specific controllers, before introducing the transformative approach—the behavior foundation model. This evolutionary trajectory not only reflects the field’s advancement toward enhanced intelligence and generalizability but also paves the way for the next-generation humanoid robot control systems.
@@ I-A Traditional Model-based Controller
Traditional WBC methods have served as the cornerstone for locomotion and manipulation in early humanoid robots [6, 7]. These methods rely heavily on physics-based models and are typically structured into a predictive-reactive hierarchy: high-level planners like centroidal model predictive control (MPC) [8, 9] generate reference trajectories, while low-level task-space whole-body controllers solve optimal control problems (OCPs) [10] to track these objectives under dynamic constraints. For instance, operational space control [11] and hierarchical task control [12] establish the theoretical foundation, while [13] advances hierarchical quadratic programming (QP) solvers and enables real-time performance in multi-task scenarios such as balance, walking, and manipulation. Whole-body operational space control (WBOSC) extends these methods by integrating the control of multiple body parts, handling redundancy, and managing dynamic constraints across the entire robot body [14, 15, 16]. These frameworks have been widely applied in humanoids, such as Atlas [4], HRP-2 [17], and DLR’s torque-controlled robots [18], achieving robust locomotion and multi-contact interactions.
Despite their success, traditional WBC systems face critical limitations: (i) task design, gain tuning, and heuristic adjustments for complex behaviors (e.g., uneven terrain or dynamic transitions) remain labor-intensive and brittle, (ii) real-time MPC struggles with high-dimensional systems, often requiring simplifications that sacrifice dynamic fidelity [19], (iii) the lack of flexibility to execute highly dynamic skills (e.g., backflips or rapid contact switches) or adapt to unforeseen disturbances [20], and (iv) weak robustness, as even a little push may topple a robot with model-based walking controller. These challenges are especially significant in humanoids, where tasks often require rich coordination, contact reasoning, and situational awareness [21, 22]. As a result, recent research increasingly shifts toward data-driven approaches, aiming to learn motor skills, coordination policies, and behavioral priors from demonstrations or reinforcement learning [23, 24].
@@ I-B Learning-based and Task-specific Controller
Learning-based methods, particularly reinforcement learning (RL) and imitation learning (IL), have emerged as promising alternatives to traditional WBC methods, enabling robots to acquire complex skills through environmental interaction or human demonstrations [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. For example, the DeepMimic framework [23] combines deep RL with motion capture data, allowing physically simulated characters to learn dynamic skills while maintaining natural motion quality. This work is further extended by adversarial motion priors (AMP) [36], which enables more stylized and diverse character control while preserving physical realism. Similarly, the HoST [37] framework introduces an RL-based standing-up controller that learns adaptive and stable motions across diverse laboratory and outdoor environments, highlighting the robustness and task-specific learning capability of RL. For IL-based approaches, the TRILL [38] approach combines virtual reality (VR) teleoperation with WBC for humanoid loco-manipulation, demonstrating an 85% success rate in real-world bimanual tasks [39]. Further advances, such as ExBody [5, 40], develop an expressive WBC framework that decouples upper-body IL (for stylistic motions) from robust lower-body locomotion, enabling humanoid robots to dynamically adapt their gait while performing diverse movements. This approach overcomes the instability of full-body imitation caused by morphological mismatches between humans and robots.
While learning-based methods have demonstrated remarkable success in diverse humanoid WBC tasks, they face fundamental challenges that limit their broader applicability. RL-based approaches suffer from sample inefficiency, often requiring millions of environment interactions to converge, while remaining highly sensitive to the reward function design—poorly shaped rewards can lead to unintended behaviors or local optima [41]. Furthermore, the simulation-to-reality (Sim2Real) gap exacerbates these limitations, as policies trained in simulation frequently degrade when confronted with real-world dynamics, sensor noise, and hardware imperfections [42, 43, 44, 45]. In contrast, IL-based methods are more sample-efficient yet pose a significant challenge to data collection, and learned policies often inherit the biases and limitations of the demonstrator [46, 47, 48, 49, 50]. Moreover, both paradigms struggle with generalization, in which learned policies typically excel only at narrow tasks and fail to adapt to new scenarios without extensive retraining. These challenges collectively underscore the need for approaches that combine the flexibility of learning with structured priors for robustness and generalizability, a gap that behavior foundation models aim to bridge.
@@ I-C Behavior Foundation Model
The term ”behavior(al) foundation model (BFM)” is first introduced in [51], which proposes a successor measure-based framework for training generalist policies capable of instantly imitating diverse behaviors from minimal demonstrations. It demonstrates that BFMs pretrained on unsupervised interaction data are promising for eliminating task-specific RL fine-tuning by solving imitation tasks through forward-backward state feature matching, while simultaneously supporting multiple IL paradigms via a unified representation, such as behavioral cloning, reward inference, and distribution matching. Subsequent work [52, 53, 54, 55, 56] has established BFMs as a class of RL agents capable of unsupervised training on reward-free transitions while yielding approximately optimal policies for broad classes of reward functions at test time without additional learning or planning.
In this paper, we extend the definition of BFMs as a specialized class of foundation models [57] designed to control agent behaviors in dynamic environments. As illustrated in Fig. 1, rooted in the principles of general foundation models (e.g., GPT-4 [58], CLIP [59], and SAM [60]) that leverage broad, self-supervised pre-training on large-scale static data, BFMs are often trained on extensive behavior data (e.g., human demonstrations, or agent-environment interactions), encoding a comprehensive spectrum of behaviors rather than specializing narrowly in single-task scenarios. This property ensures that the model can readily generalize across different tasks, contexts, or environments, demonstrating versatile and adaptive behavior generation capabilities. Recent advancements in vision-language-action (VLA) models [59, 61, 62, 63] have focused on integrating vision, language, and action to handle multi-modal tasks, excelling in dynamic settings where they generate context-aware responses based on visual and linguistic inputs. In contrast, BFMs are primarily designed for directly controlling agent behaviors such as locomotion, manipulation, and interaction. Moreover, most existing VLA models are applicable to relatively stable platforms, such as mechanical arms or wheeled humanoid robots [64], whereas BFMs are designed to handle the sophisticated WBC of full-size humanoid robots.
@@ II-A Humanoid Whole-body Control
Humanoid robots are expected to operate in unstructured environments that are dynamic and unpredictable, demanding control systems that are highly versatile, robust, reconfigurable, dexterous, and mobile compared to less agile robotic platforms [65]. To that end, the humanoid WBC is proposed to coordinate the motion of multiple robots’ appendages to execute multiple tasks simultaneously and reliably. It considers the entire robot body as a single and integrated system, managing locomotion, manipulation, and interaction with the environment using a unified set of control algorithms [14]. As illustrated in Fig. 2, the humanoid WBC has evolved from traditional model-based approaches to flexible learning-based approaches, moving toward a generalist that solves broad tasks in diverse scenarios [65, 64, 5, 66, 67, 68]. In line with this trend, BFMs have emerged as a promising approach to achieve general-purpose WBC through large-scale pre-training on diverse motion data, which will be comprehensively discussed in the following contents.
@@ II-B Reinforcement Learning
We frame the RL problem as a Markov decision process (MDP), which can be defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r^{\rm ext},\gamma)$ [69]. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(\bm{s}^{\prime}|\bm{s},\bm{a})$ is the probability measure on $\mathcal{S}$ defining the stochastic transition to the next state $\bm{s}^{\prime}$ obtained by taking action $\bm{a}$ in state $\bm{s}$, $r^{\rm ext}(\bm{s}):\mathcal{S}\rightarrow\mathbb{R}$ is the extrinsic/task reward function, and $\gamma\in[0,1]$ is a discount factor. A policy $\pi$ is defined as the probability measure $\pi(\bm{a}|\bm{s})$ that maps each state to a distribution over actions. Furthermore, we denote $\mathrm{Pr}(\cdot|\bm{s}_{0},\bm{a}_{0},\pi)$ and $\mathbb{E}(\cdot|\bm{s}_{0},\bm{a}_{0},\pi)$ as the probability and expectation operators under state-action sequences $(\bm{s}_{t},\bm{a}_{t})_{t\geq 0}$ starting at $(\bm{s}_{0},\bm{a}_{0})$ and following the policy $\pi$ with $\bm{s}_{t}\sim P(\bm{s}_{t}|\bm{s}_{t-1},\bm{a}_{t-1})$ and $\bm{a}_{t}\sim\pi(\bm{a}_{t}|\bm{s}_{t})$. The goal of RL is to learn a policy that maximizes the expected discounted return:
@@ III-A1 Goal-conditioned Learning
As shown in Figure 4, goal-conditioned learning is a framework in RL where an agent’s behavior is conditioned on a specific goal or objective, typically provided as input. Unlike traditional RL, where the agent learns from raw state-action pairs without explicit task-specific guidance, goal-conditioned learning integrates the goal into the agent’s policy, enabling it to adapt its actions toward achieving that specific goal. The goal can be specified in various forms, such as a target state, an objective function, or an external task description [70, 71, 72]. The key advantage of goal-conditioned learning lies in its ability to learn a more flexible and transferable policy that can be applied to a wide range of tasks, as it directly incorporates the task’s goal during training, rather than requiring retraining for each specific task. This makes it particularly useful in environments where the agent needs to solve multiple tasks or interact with changing environments.
@@
TeamPlay [73], ASE [74], PHC [70], CALM [75], MTM [71],
CASE [76], InterMimic [77], MoConVQ [78], H-GAP [72],
MaskedMimic [79], HugWBC [66], HOVER [67], ModSkill [80],
CLONE [81], TWIST [82], TWIST2 [83], AMS [84],
Any2Track [85], BFM4Humanoid [86], SONIC [87]
@@
ICM [88], DIAYN [89], RND [90], APS [91],
ProtoRL [92], RE3 [93], ODPP [94]
@@
FB [95], FB-IL [51], FB-AW [52], FB-ARE [52],
FB-AWARE [52], Motivo [53], BFM-Zero [96]
Adaptation Adaptation
@@
TokenHSI [97], Belief-FB [56], Rotation-FB [56],
TaskTokens [55], ReLA [54], LoLA [54]
@@
UniHSI [98], TokenHSI [97], UniPhys [99], CLoSD [100],
LangWBC [101], LeVERB [102], BeyondMimic [103], SENTINEL [104]
Skill learning from motion tracking. Among the diverse approaches to goal-conditioned learning, tracking-based learning represents a specialized form where the target behavior is explicitly defined by dense reference supervision or guidance, typically derived from motion capture data or expert demonstrations. At each time step, the agent is often trained to track the given reference motion’s joint angle or kinematic pose of the next time step [23]. The primary motivation behind tracking-based learning is that learning to track a single pose is more achievable and general than directly imitating a whole motion, especially a complex motion.
For example, the TeamPlay [73] framework trains an agent to imitate a large amount of football motion capture data via a DeepMimic-like [23] approach, aiming to learn a complete behavior set required for the football game. Then the agent is leveraged to sample substantial state-action pairs to train a neural probabilistic motor primitive (NPMP) model [105] and derive a low-level latent-conditioned controller. Finally, the controller is applied for further drill learning by conducting RL with a drill-specific (e.g., follow, dribble, shoot, and kick-to-target) reward function. Here, the learned low-level controller can be viewed as a BFM as it learns realistic human-like movements based on motion capture data and can be rapidly adapted to diverse higher-level drill learning.
Similarly, the adversarial skill embeddings (ASE) framework [74] learns a reusable latent space of motor skills by combining adversarial IL with unsupervised RL. Trained on unstructured motion data, ASE produces a latent-conditioned low-level controller capable of generating diverse and physically plausible behaviors, serving as a general-purpose motor prior to downstream tasks. Building on ASE, the conditional adversarial latent models (CALM) framework [75] incorporates a conditional discriminator to enable fine-grained control over generated motions through latent manipulation. This line of work is further extended by CASE [76], which introduces skill-conditioned IL with training techniques such as focal skill sampling and skeletal residual forces to enhance agility and motion diversity.
@@
While the methods above achieve efficient skill acquisition from large behavior datasets, the HugWBC [66] method explores learning versatile locomotion skills without relying on pre-collected motion data. The framework automatically generates adaptive behaviors through a structured RL process, where a general command space dynamically produces feasible velocity, gait, and posture targets during training. By reformulating WBC as a self-supervised command-tracking problem, the work establishes a new direction for developing general-purpose humanoid controllers that learn robust skills through environmental interaction rather than data imitation.
@@
Moving beyond body-level skill learning, ModSkill [80] introduces a modular framework that decouples full-body motion into part-specific skills for individual body parts. This modularization allows for efficient and scalable learning, as each body part is controlled independently by a low-level controller driven by part-specific skill embeddings. ModSkill’s ability to focus on body-part-level skills makes it a powerful system for controlling complex motions and adapting learned behaviors across different tasks. By utilizing a skill modularization attention layer, ModSkill enhances the generalization of motor skills across various tasks, such as reaching or striking, thereby further improving task-specific adaptation.
@@
From primitive skills to high-level goal execution. The success of BFMs in learning diverse primitive skills has propelled the development of more advanced BFMs capable of interpreting and executing high-level goals, including language instructions and multi-task objectives. A notable example is MoConVQ [78], which introduces a unified motion control framework based on discrete latent codes learned via vector quantized variational autoencoder (VQ-VAE). The model supports a wide range of downstream tasks, including motion tracking, interactive control, and text-to-motion generation, by offering a compact and modular representation. MoConVQ also integrates with large language models (LLMs) and enables the simulated agents to be directed via in-context language prompts, thereby bridging symbolic reasoning and physical control.
@@
Figure 5: MaskedMimic [79] is a representative BFM trained via goal-conditioned learning, which supports diverse control modalities and achieves seamless transitions between distinct tasks.
@@
Meanwhile, MaskedMimic [79] addresses physics-based character control as a general motion inpainting problem, producing full-body motions from partial descriptio, such as masked keyframes, objects, or text instructions. MaskedMimic involves a two-phase training process. First, a fully-constrained motion tracking controller is trained to imitate diverse reference motions. Then, a partially-constrained VAE-based policy distills this knowledge through masked goal conditioning. As a result, MaskedMimic can dynamically adapt to complex scenes and support applications ranging from VR control to complex human-object interaction (HOI), as illustrated in Fig. 5. In addition, InterMimic [77] focuses on the HOI scenario and designs a two-stage teacher-student framework that distills imperfect motion capture interaction data into robust and physics-based controllers. Teacher policies are trained on subsets of noisy data and refined through simulation, then distilled into a student policy with RL-based fine-tuning. This curriculum strategy enables generalization across diverse interactions with high physical fidelity.
@@
For real-world robotic applications, HOVER [67] introduces a multi-mode policy distillation framework that enables humanoid robots to switch seamlessly between tasks (e.g., locomotion, manipulation, and navigation) using a single unified policy distilled from an oracle. This eliminates the need for task-specific controllers, demonstrating general-purpose control in real-world environments, akin to the versatility seen in MaskedMimic for virtual characters.